OpenStudy (anonymous):

What is the simplified form of 3 over 2x minus 5 + 21 over 8 x squared minus 14x minus 15 ? 6 times the quantity 2 x plus 5 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 times the quantity x plus 1 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 time the quantity x plus 1 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 times the quantity x plus 1 end quantity over the quantity 2x plus 5 end quantity times 4 x plus 3

OpenStudy (anonymous):

@Baby_Bear69

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

@♪Chibiterasu

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

ello love :D need some help? @Motown117

OpenStudy (anonymous):

yes i need help

Nnesha (nnesha):

\[\huge\rm \frac{ 3 }{ 2x-5 } +\frac{21}{8x^2-14x-15}\] like this ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

first factor the quadratic equation

OpenStudy (anonymous):

can you walk me through how to do this?

OpenStudy (anonymous):

you got this! :D @nnesha

Nnesha (nnesha):

alright \[\huge\rm 8x^2-14x-15\] multiply AC (a =leading coefficient, c=constant term) ,

Nnesha (nnesha):

8 times -15 = -120 now find two number when we multiply them we should get product of AC and when we add or subtract them we should get middle term which is -14 in this equation

OpenStudy (anonymous):

7times -7

OpenStudy (anonymous):

thats wrong isent it?

Nnesha (nnesha):

hmm this is AC method to factor the product of 2 number should be equal to AC and sum = -14 so find two number if you multiply them you should get -120( product of a and c =8 times -15)

Nnesha (nnesha):

and when you add/subtract them you should get -14

OpenStudy (anonymous):

6 times -20

Nnesha (nnesha):

right and since the leading coefficient isn't one we need to factor by grouping  carry down the first and last term and write both numbers in middle